Trigonometry

what is cos90

riya garg
11 Years agoGrade

3 Answers

Jitender Singh
11 Years ago
Ans:
Hello Student,
Please find answer to your question below
0, 30, 45, 60 & 90 degrees comes under the standard angles
cos90^{0} = 0
cos0^{0} = 1
cos30^{0} = \frac{\sqrt{3}}{2}
cos60^{0} = \frac{1}{2}
Harshakavya
11 Years ago
cos(90) = 0
Nidhi Prajapati
11 Years ago
Cos 90 ie.cosπ/2= 0